OpenStudy (anonymous):

Find a complex Z such that : |Z|+Z=4(2+i).

OpenStudy (amistre64):

since |Z| = the distance from the origin id propose something like: \[\sqrt{a^2+b^2}+a+bi=8+4i\]

OpenStudy (amistre64):

\[\sqrt{a^2+b^2}+a+bi=8+4i\] \[\sqrt{a^2+b^2}=8+4i-a-bi\] \[a^2+b^2=(8+4i-a-bi)^2\] \[a^2+b^2=a^2+2abi-(16+8 i) a-b^2+(8-16i) b+(48+64i)\] \[b^2=2abi-(16+8 i) a-b^2+(8-16i) b+(48+64i)\] \[0=2abi-(16+8 i) a-2b^2+(8-16i) b+(48+64i)\] that might be going someplace, but it doesnt seem like it

OpenStudy (amistre64):

maybe: \[|Z|=8+4i=\sqrt{8^2+4^2}\] \[|Z|=\sqrt{80}\] \[|Z|=4\sqrt{5}\] \[4\sqrt{5}+Z=8+4i\] \[Z=(8-4\sqrt{5})+4i\] maybe?

OpenStudy (amistre64):

z=3+4i according to the wolf

OpenStudy (anonymous):

\[Z = a + ib\] \[ \sqrt{a^2 + b^2} + a + ib = 8 + i4\] \[b=4\] \[\sqrt{a^2 + b^2} + a = 8 \] \[\sqrt{a^2 + 16} = 8 -a\] \[a^2 + 16 = 64 + a^2 - 16a\] Maybe like this I am not sure

OpenStudy (anonymous):

"according to the the wolf" LOL

OpenStudy (amistre64):

seems like (|Z|+a) = 8

OpenStudy (anonymous):

\[16a = 64 - 16\] \[4a = 16 - 4\] \[a = 4-1\] \[a = 3\] \[a + ib = z= 3 + i4\]

OpenStudy (anonymous):

Yeah exactly!

OpenStudy (amistre64):

comparing coeffs, good idea

OpenStudy (phi):

In case it is not obvious |Z| is by definition a real number. This means you can break the problem into 2 parts: let Z= a+bi Im(Z)= Im(8+4i)= 4 so b= 4 and Re(|Z|+Z)= Re(8+4i)= 8 |Z| + a = 8 or |Z|= 8-a substitute for |Z|= sqrt(a^2+b^2) = sqrt(a^2+16), and solve for a As noted in previous posts, a= 3

OpenStudy (anonymous):

Solution : Z=3+4i.

OpenStudy (phi):

And note it is easily checked. from pythagoras |Z|= 5 and 5+3+4i= 8+4i